Three resistors connected in parallel have individual values of 6.0, 10.0 and 12.0 Ohm, respectively. If this combination is connected in series with a 15-V battery and a 3.0 Ohm resistor, what is the current in the 12 Ohm resistor?
Answer #1

Let us find the equivalent resistance of 3 resistors in parallel

1/RP = 1/R1 + 1/R2 + 1/R3

1/RP = 1/6 + 1/10 + 1/12

RP = 2.86 ohm

 

I net = V/ (R+RP)

= 15/(3+2.86)

= 2.56 A

 

V across 3 ohm = I*3 = 2.56*3 = 7.68 V

 

V across parallel component = 15 - 7.68 = 7.32 V

 

I in 12 ohm resistor = V across 12 ohm / 12 ohm

= 7.32/12

= 0.61 A

Answer: 0.61 A